OpenStudy (anonymous):

Find the x and y intercepts. Please show all work. -4x+5y=-20

OpenStudy (angela210793):

1ST TAKE X=0 5Y=-20 y=-4 (0;-4) then take y=0 -4x=-20 x=5 (5;0)